### v1

# Through the Looking Glass: Bitcoin Treasury Companies

Abstract
‘Bitcoin treasury companies’ have taken stock markets by storm amassing billions of dollars worth of tokens. More than a hundred companies active in various crypto-tokens, not just Bitcoin, are listed. The ability to generate leverage, for example stemming from borrowing collateralised by stock portfolios, helps to understand this phenomenon.
In addition, the extension of the binary-choice Kelly criterion to incorporate uncertainty in the form of the Kullback-Leibler divergence or more generally Bregman divergence is studied.
I Introduction
The Bitcoin whitepaperbtc dates back to October 2008, the height of the ‘great financial crisis’ (GFC) and only about a month after the Lehman bankruptcy. Legacy finance, as it is now sometimes called, was in disarray and a search for alternatives gathered pace. Bitcoin, maybe initially more of a Gedankenexperiment than a practical means of exchange, has over the intervening years grown dramatically. The associated coin BTC now has a market capitalisation of more than two trillion. Thousands of other coins have joined the proverbial gold rush. For years legacy finance largely ignored these upstarts, but the growth of BTC’s and other tokens’ market capitalisations has piqued curiosity and greed.
Companies, and not just individuals, have started hoarding bitcoins. There are currently over a hundred companies that are devoted to this one task - the holding of crypto-assets. Many market themselves as vehicles whose mission it is to acquire ever-more and ever-faster crypto-tokens financed by a mixture of equity and various form of debt issuance. Puzzlingly, their market capitalisation can be significantly higher than the net asset value of their crypto-assets. This is especially surprising when the non-crypto business of these companies is often between insignificant and non-existent. How can this be? The paper suggests a possible answer, which goes beyond irrational exuberance.
In the next section popular new performance indicators are briefly introduced and in the subsequent section the role of leverage is explored. The penultimate section of the paper is devoted to an extension of the one-dimensional Kelly criterion to cases where the winning probability is uncertain. A conclusion rounds of the paper.
II Key performance indicators ‘reimagined’
One of the key performance indicators (KPIs) touted by crypto-treasury companies is mNAV, which is the ratio of enterprise value [^1] In the case of an ‘idealised’ company, which is the only company we consider, the enterprise value is equivalent to the market capitalisation. We assume the company is debt free, no preferred equity has been issued and (for completeness sake) no pension liabilities exist. to the value of the crypto-tokens held. If this is above one issuing equity to purchase further tokens is value accretive. To see this take a simple example. Assume the token trades at \$$\$1$, and the mNAV is $4$. Further assume $4$ shares have been issued trading at \$$\$1$ each, and one token is in the companies treasury. Next a share is issued at the prevailing price of \$$\$1$ and one token is bought. Assuming an unchanged mNAV, the price per share becomes \$ $\$\,8/5$ an increase of $80\%$. Magically, everyone seems to be better off. How can this be? Is it due to a suspension of disbelief, since the gap between net asset value (NAV) and the market capitalisation has widened [^2] There are also tax and other hurdles that prevent investors from buying tokens directly. For example, in Japan holding Bitcoin can be less tax efficient than stock ownership. Cryptocurrency capital gains tax in Japan is up to $55\%$, while the comparable tax on stocks is only about $20\%$. These tax rates are currently under review.. Why overpay for fractional ownership of crypto-tokens via an intermediary, when either through direct ownership or some other intermediary, as should be possible in a competitive market, the same exposure can be obtained cheaper?
‘Bitcoin per share’, the fraction of Bitcoin held per share, is also a popular KPI. A rise ‘Bitcoin per share’ is viewed as beneficial for the shareholders. This can be achieved with debt or with equity issuance, as we saw above, for mNAV larger than one. Debt issuance is interesting, since it enables the shift of risk to outside parties. This is also true for borrowing by shareholders against portfolio value. The paper considers in the next section credit transfers by borrowing of shareholders but not by the company itself. The different types of credit transfers are not interchangeable, even if in both cases risk is shifted away from the ‘centre’.
Other KPIs including BTC yield, BTC gain and BTC \$ gain have also been bandied about but are maybe not as helpful, or at least add little beyond the metrics mentioned above, to understand the price behaviour. Once mNAV approaches one, other KPIs as well as price impact [^3] For an analysis of strategies incorporating price impact see bkm2016 . Here we are interested with the price impact on Bitcoin by companies buying a substantial amount of the outstanding tokens. gain in importance. In the next section we return to mNAV.
III Mirror images: Is it a virtuous, a vicious or a virtual cycle?
In section II of a paper on ‘rational market failures’ meister2022 a cyclical strategy was described that allowed risk to be passed from stock investors to credit providers. With sufficient market impact an investor was able to extract money. Next, an outline of how the process works. The investor would buy a particular stock, drive up the stock price, since one-sided demand can lift the price, and the portfolio value. These portfolio gains the investor could use as collateral to borrow against. This would lead to a virtuous cycle for the stockholders, and a potentially vicious cycle for the lenders, which would accumulate risk. In extreme cases the investor could extract cash and come out ahead, even if the stock eventually crashes. For details, see the earlier paper.
The ‘innovation’ of crypto-treasury companies is to exploit the counter-intuitive situation of having a mNAV larger than one. In its simplest form the issuance of stock converted back into crypto-tokens raises the stock price, followed by more stock issuance. Additional stock in one scenario is bought by existing holders of the stock, who borrow against the ever increasing value of their stock portfolio to finance the new purchases. This is true even without any overt price impact on crypto-tokens. It wouldn’t be surprising if the steady buying by treasury companies of an assets that is in the case of Bitcoin of fixed supply will in addition raise the price of the token. The efficiency and thereby viability of the equity-issuance strategy only depends on banks or other credit providers treating stock of Bitcoin treasury companies like other assets that can be deposited as collateral for loans with a reasonable haircut. When collateral requirements are raised, the attractiveness of these companies as endlessly equity-issuing financial perpetuum mobiles may disappear suddenly.
IV The Kelly criterion in uncertain markets
The one-dimensional Kelly criterion is in this section extended to uncertain winning probabilities. The Kullback-Leibler (KL) distance or some more general Bregman divergence between the assumed and real probabilities captures the idea of lack of knowledge. The application of the KL distance to the Kelly criterion was already discussed in meister2024 but in the earlier paper the emphasis was on large deviations and the ability to distinguish strategies associated with different investment fractions.
Before a toy model for uncertainty and the influence on the investment fraction is discussed, we study the deliberate over- or under-investment in a situation with a known winning probability. It was shown in meister2011 how this affects the portfolio growth rate negatively as well as asymmetrically. This result can be understood intuitively, since if one underinvests, the growth underperforms but is bounded by zero. If one instead over-invests, the growth is not bounded by zero and instead can turn negative. Next, the intuition is confirmed by expanding the growth rate in terms of a small parameter linked to the optimal fraction. If the winning probability is known to be $p$ in the standard double-or-nothing game, and one varies the optimal investment fraction by an infinitesimal amount $\epsilon$, i.e. $f=2p-1+\epsilon$, then this leads to
$p\log(1+f)+(1-p)\log(1-f)=p\log(2p+\epsilon)+(1-p)\log(2-2p-\epsilon)$
$=$ $H(p)+1+p\log\Big{(}1+\frac{\epsilon}{2p}\Big{)}+(1-p)\log\Big{(}1-\frac{\epsilon}{2-2p}\Big{)}$
$=$ $H(p)+1+p\Big{(}\frac{\epsilon}{2p}-\frac{\epsilon^{2}}{8p^{2}}+\frac{\epsilon^{3}}{24p^{3}}\Big{)}-(1-p)\Big{(}\frac{\epsilon}{2-2p}+\frac{\epsilon^{2}}{8(1-p)^{2}}+\frac{\epsilon^{3}}{24(1-p)^{3}}\Big{)}+O\big{(}\epsilon^{4}\big{)}$
$=$ $H(p)+1-\frac{1}{8(1-p)p}\epsilon^{2}-\frac{2p-1}{24(1-p)^{2}p^{2}}\epsilon^{3}+O\big{(}\epsilon^{4}\big{)}.$
Let us phrase the above result in terms of the existence and sign of the quadratic and cubic terms in $\epsilon$. The growth is suboptimal for non-zero $\epsilon$, i.e. the leading term is quadratic and has a negative sign, and the loss in the portfolio growth rate due to over-investment is larger than due to under-investment for the same absolute value in $\epsilon$, i.e. the cubic term has a negative sign.
The inputs of the toy model for investment uncertainty discussed next are the value of the KL divergence $D_{\mathrm{KL}}\left(q\parallel p\right)$ set equal to $\alpha$, which is assumed to be small, i.e. $\alpha\ll 1$, as well as the probability $q$, which is assumed to be larger than $1/2$. The separation of $q$ from the unknown ‘correct’ probability is given by the KL divergence $\alpha$. To avoid possible solutions for $p$ from falling below $1/2$, which would cause the associated investment fraction to be negative, $q$ is set sufficiently above $1/2$. We further define $p=q+\epsilon$, where $\epsilon$ is small. An expansion of $D_{\mathrm{KL}}\left(q\parallel p\right)$ in terms of $\epsilon$ has the form
$D_{\mathrm{KL}}\left(q\parallel p\right)=p\log\Big{(}\frac{p}{q}\Big{)}+(1-p)\log\Big{(}\frac{1-p}{1-q}\Big{)}$
$=$ $\sum_{n=1}^{\infty}\Big{(}q+\epsilon\Big{)}(-1)^{n}\frac{\epsilon^{n}}{n\,\,q^{n}}-\Big{(}1-q-\epsilon\Big{)}\frac{\epsilon^{n}}{n\,\,(1-q)^{n}}$
$=$ $\sum_{n=2}^{\infty}\frac{\epsilon^{n}}{n^{2}-n}\Big{(}\frac{(-1)^{n}}{q^{n-1}}-\frac{1}{(1-q)^{n-1}}\Big{)}=\sum_{n=2}^{\infty}\frac{\epsilon^{n}}{n^{2}-n}\Big{(}\frac{(-1)^{n}(1-q)^{n-1}-q^{n-1}}{q^{n-1}(1-q)^{n-1}}\Big{)}$
$=$ $\frac{1}{2(1-q)q}\epsilon^{2}-\frac{1}{6}\frac{(1-q)^{2}+q^{2}}{(1-q)^{2}q^{2}}\epsilon^{3}+\frac{1}{12}\frac{(1-q)^{3}-q^{3}}{(1-q)^{3}q^{3}}\epsilon^{4}-\frac{1}{20}\frac{(1-q)^{4}+q^{4}}{(1-q)^{4}q^{4}}\epsilon^{5}+O(\epsilon^{6}).$
If, as stated above, both $D_{\mathrm{KL}}\left(q\parallel p\right)$, equal to $\alpha$, and $q$ are fixed, then in lowest approximation, as one solves the equation up to a quadratic terms, one gets two solutions for $\epsilon$. Various cases can be considered. One can consider the worst case, i.e. the minimal solution for $p$, the best case, i.e. the maximal solution for $p$, and a ‘least informative’ case, i.e. with equal weights for both possible solutions of $p$. In the equal weighted case, due to the asymmetry associated with over- and under-investment, as described above, the optimal investment fraction is lower than $12q-$ 1. To summarise, uncertainty about the right value for the winning probability leads to a lowering of the growth rate and the optimal fraction. An attractive formula for the adjusted investment fraction that captures this heuristically is $f^{*}\exp(-\lambda\alpha)$, where $f^{*}$ is the optimal uncertainty-free fraction $12q-$ 1, and $\lambda$ is a risk-aversion parameter. This result can be extended from KL divergence to the more general Bregman divergence by doing the relevant Taylor expansions for cases of interest like the Itakura–Saito distance. A similar ‘conservatism’ can be observed.
The conclusion attempts again to shed light on treasury companies.
V Conclusion
Crypto treasury companies with a rising ‘Bitcoin per share’ and mNAV above one seem to have discovered a financial perpetuum mobile and are eager to recruit followers, since copy-cats compete for the restricted supply of Bitcoins and in the process push up prices. Is there a fault in their logic? First, credit risk aggregates in the system, when these stocks are used as collateral to obtain leverage to race ahead in the accumulation process. This risk has to be managed to mitigate potential losses, when there is a dramatic price reversal of BTC or mNAV sufficiently compresses leading to portfolio liquidations. Second, if price impact of buying limited supply tokens is central to the scheme, then there is the possibility of a widening gap between the price and the value of these tokens. As Fischer Black once stated for equities, the gap between price and value in an ‘efficient market’ [^4] Fischer Black in “Noise”, Journal of Finance, vol. 41, pp. 529–543 (1986) stated that he thought ‘markets are efficient almost all of the time. “Almost all” means $90\%$ of the time’. This statement was and to a lesser extent still is controversial. is ‘almost all’ of the time not larger than $100\%$. Constraints on value exists even for crypto tokens, since a market capitalisation much beyond the current single digit trillion dollar amount has a good chance to unbalance not just the financial system and the economy in general, but also politics.
Can the gap between market capitalisation and net asset value hold indefinitely as more equity is issued, bought by current holders, and used as collateral for loans to buy further tokens? This seems unlikely, since the strategy has a weak point: Credit providers. The willingness of credit providers to lend against stocks is surely limited, since their risk appetite is bounded. What can credit providers do? They should be cognizant of the gap between market capitalisation and NAV and align their lending to the limited liquidation value under market stress of large crypto-token portfolios. It seems brokers are aware of the risk and are restricting lending against concentrated positions in token treasury companies [^5] This is suggested by Reddit discussion forums and similar anecdotal evidence.. This could break the cycle, but, as the saying goes, ‘the market can stay irrational longer than investors can stay solvent’.
The history of finance is full of wondrous schemes. After the South Sea bubble even Newton was supposed to have said, ‘I can calculate the motion of heavenly bodies, but not the madness of people’, and similar events in France, where the ‘Mississippi Bubble’ ended badly for the buccaneer John Law, undermined the established order, and prepared the ground for the coming revolution.
In the second part of the paper, simple information theoretic ideas are employed to understand the influence of uncertainty on decision making, and an asymmetric effect on the optimal fraction for the binary-choice Kelly criterion was observed as the winning probability becomes uncertain.
The paradox of high mNAV crypto treasury companies is captured by the White Queen in Lewis Carroll’s ‘Through the Looking Glass’, when she says, ‘[t]he rule is, jam to-morrow and jam yesterday – but never jam to-day’.
  1. (1) J. Kelly, A new interpretation of information rate. Bell System Technical Journal 35(4): 917–926 (1956).
  2. (2) Y. Lv & B. K. Meister, “Applications of the Kelly Criterion to multi-dimensional Diffusion Processes” International Journal of Theoretical and Applied Finance 13, 93-112. (2010) 27, reprinted in ”The Kelly Criterion: Theory & Practice:”, ed. by MacClean, Thorpe & Ziemba, World Scientific. 285-300. (2011).
  3. (3) B. K. Meister, “Meta-CTA Trading Strategies based on the Kelly Criterion”, arXiv:1610.10029. (2016).
  4. (4) B. K. Meister, “Meta-CTA Trading Strategies and Rational Market Failures”, arXiv:2209.05360. (2022).
  5. (5) B. K. Meister, “Gambling the World Away: Myopic Investors”, arXiv:2302.13994. (2023).
  6. (6) B. K. Meister, “Application of the Kelly Criterion to Prediction Markets ”, arXiv:2412.14144. (2024).
  7. (7) S. Nakamoto, ‘Bitcoin: A Peer-to-Peer Electronic Cash System’ (2008).

### v2

@@
 Abstract
-‘Bitcoin treasury companies’ have taken stock markets by storm amassing billions of dollars worth of tokens.
-More than a hundred companies active in various crypto-tokens, not just Bitcoin, are listed.
-The ability to generate leverage, for example stemming from borrowing collateralised by stock portfolios, helps to understand this phenomenon.
-In addition, the extension of the binary-choice Kelly criterion to incorporate uncertainty in the form of the Kullback-Leibler divergence or more generally Bregman divergence is studied.
+Bitcoin treasury companies have taken stock markets by storm amassing billions of dollars worth of tokens in hundreds of entities.
+The paper discusses, how leverage - whether created through corporate debt or investors using stock as loan collateral - fuels this trend.
+The extension of the binary-choice Kelly criterion to incorporate uncertainty in the form of the Kullback-Leibler divergence or more generally Bregman divergence is also briefly discussed.
 I Introduction
@@
 How can this be?
-The paper suggests a possible answer, which goes beyond irrational exuberance.
-In the next section popular new performance indicators are briefly introduced and in the subsequent section the role of leverage is explored.
+The paper suggests possible answers, which go beyond irrational exuberance.
+In the next section popular new performance indicators are briefly introduced and in the subsequent sections the role of leverage is explored for both investors and the treasury companies themselves.
 The penultimate section of the paper is devoted to an extension of the one-dimensional Kelly criterion to cases where the winning probability is uncertain.
@@
 II Key performance indicators ‘reimagined’
-One of the key performance indicators (KPIs) touted by crypto-treasury companies is mNAV, which is the ratio of enterprise value [^1] In the case of an ‘idealised’ company, which is the only company we consider, the enterprise value is equivalent to the market capitalisation.
+One of the key performance indicators (KPIs) touted by crypto-treasury companies is mNAV, which is the ratio of enterprise value [^1] In the case of an ‘idealised’ company, which is the only company we consider in this section, the enterprise value is equivalent to the market capitalisation.
 We assume the company is debt free, no preferred equity has been issued and (for completeness sake) no pension liabilities exist.
@@
 To see this take a simple example.
-Assume the token trades at \$$\$1$, and the mNAV is $4$.
-Further assume $4$ shares have been issued trading at \$$\$1$ each, and one token is in the companies treasury.
-Next a share is issued at the prevailing price of \$$\$1$ and one token is bought.
-Assuming an unchanged mNAV, the price per share becomes \$ $\$\,8/5$ an increase of $80\%$.
+Assume the token trades at \$1\\$1, and the mNAV is $4$.
+Further assume $4$ shares have been issued trading at \$1\\$1 each, and one token is in the companies treasury.
+Next a share is issued at the prevailing price of \$1\\$1 and one token is bought.
+Assuming an unchanged mNAV, the price per share becomes \$ 8/5\\$\,8/5 an increase of $80\%$.
 Magically, everyone seems to be better off.
@@
 In the next section we return to mNAV.
-III Mirror images: Is it a virtuous, a vicious or a virtual cycle?
+III Mirror images: Is collaterlized borrowing a virtuous, a vicious or a virtual cycle?
 In section II of a paper on ‘rational market failures’ meister2022 a cyclical strategy was described that allowed risk to be passed from stock investors to credit providers.
@@
 When collateral requirements are raised, the attractiveness of these companies as endlessly equity-issuing financial perpetuum mobiles may disappear suddenly.
-IV The Kelly criterion in uncertain markets
+Next a look at companies instead of investors raising leverage through debt.
+IV Is Debt the Fulcrum of Finance Treasury Companies?
+This section considers a Bitcoin treasury company raising money through a zero coupon bond to add to an existing trove of tokens.
+This is a simplification of what one observes in markets, since treasury companies are often serial issuers with an eclectic mix of offerings.
+Nevertheless, this pared down model provides insights.
+Next, the relevant quantities are defined.
+The Bitcoin price is $B_{0}$ at time $t=0$ and described by a geometric Brownian motion with volatility $\sigma$ and drift $\mu$,
+$dB_{t}=\mu B_{t}dt+\sigma B_{t}dW.$
+The quantities A&YA\,\&\,\,Y are denominated in $[BTC]$, $C_{BS}\,\,\&\,\,P_{BS}$ in [\$/BTC][\\$/BTC] [^4] Option prices are often quoted dollars, and the fact that the price is per contract (or certain number of shares) is implicit., $X$ in [\$][\\$], $T$ in $^{\prime}time^{\prime}]$, $B_{t}$ in [\$/BTC][\\$/BTC], r&Rr\,\,\&\,\,R in $1/^{\prime}time^{\prime}]$, $\sigma^{2}$ in $^{\prime}time^{\prime}]$ and $Q$ is a dimensionless number.
+Pre-issuance the amount of Bitcoin held in the company is $A$.
+At time $t=0$, $Q$ units of a zero-coupon bond are issued at price $X$ per unit and are to be redeemed by the company for 1\$1\\$ per unit at time $T$.
+The issuance proceeds of $XQ$ are immediately converted into $XQ/B_{0}$ Bitcoins.
+The company holds a fixed amount of Bitcoins, i.e.
+$A+XQ/B_{0}:=Y$, until the bonds mature.
+Bankruptcy can only occur at $T$.
+At $T$ the following happens, the bond investors receive
+$Min\Big{[}1\$\,Q,B_{T}Y\Big{]}$
+and the residual or unwind value of the company is
+$Max\Big{[}B_{T}Y-1\$\,Q\,,0\Big{]}.$
+We take $B_{0}\,\,\,\&\,A$ to be inputs, whereas the amount $Q$ is to be optimised to maximise the post-issuance share price, i.e. just after $t=0$.
+What is the value of the stock post debt issuance?
+Since the company has an obligation to service its the debt, the stock is an option on Bitcoin.
+The value at time $t=0$ is given by $Y$ units of the Black-Scholes formula for an European call
+$Y\,\,C_{BS}(B_{0},1\$\,Q/Y,\sigma,R,r,T),$
+with $B_{0}$ the price of the underlying, 1\$Q/Y1\\$\,Q/Y the strike [^5] Here and elsewhere 1\$1\\$ is added, to have dimensional consistency., $\sigma$ the Bitcoin implied volatility, $T$ the maturity, and for completness R&rR\,\,\&\,\,r are the relevant interest rates.
+The Bitcoin yield curve was modelled theoretically in bhm and empirically in new provides $r$, while $R$ is the relevant interest rate for the riskless USD curve.
+The Bitcoin interest rate applied in the Black-Scholes formula can be modified by storage cost and convenience yield.
+For short maturities the price can also be interpolated from data found on Deribit or other exchanges.
+If the Bitcoin to dollar exchange rate drops below $Y/(1\$Q)\geq B_{T}$ the option expires worthless and the company retains zero value.
+In this, like the rest of the analysis, bankruptcy cost is ignored.
+Similarly, one can also determine the value of $X$, since the package of an appropriately structured put and the risky issued bond is a risk-free bond.
+In more detail, $XQ$ is the difference between the risk-less price, which only depends on discounting, i.e.
+$1\$\,e^{-RT}Q$, and an put embedded in the bond, since the $Y$ Bitcoins held by the company will be insufficient the repay the obligations at $T$, if the Bitcoin prices falls below $B_{T}\leq 1\$Q/Y$.
+$X$ is calculated using again the Black-Scholes model, or based on prices observed on the cryptocurrency options exchange Deribit or similar venues.
+The expression $(1\$\,e^{-RT}-X)Q$ is equivalent to the Black-Scholes put price for $Y$ units of $P_{BS}(B_{0},1\$\,Q/Y,\sigma,R,r,T)$, since the put payout has the form
+$Max\Big{[}1\$\,Q\,-B_{T}Y,0\Big{]},$
+and compensates for the shortfall the bond investors suffer, if the company goes bankrupt.
+Put and call prices are linked by put-call parity, i.e.
+$YC_{BS}(B_{0},1\$\,Q/Y,\sigma,R,r,T)-YP_{BS}(B_{0},1\$\,Q/Y,\sigma,R,r,T)=YB_{0}e^{-rT}-1\$\,Qe^{-RT}$.
+As stated above, one can view the share of the treasury company for $0<t\leq T$ as an option on the value of Bitcoin.
+If $Y\,C_{BS}(B_{0},1\$\,Q/Y,\sigma,R,r,T)$ for a particular choice of $Q$, or alternatively the price for this structure on Deribit, is larger than $B_{0}A$, then issuing bonds creates instantaneous value for shareholders, and an instability exists until the issuance feedback loop is exhausted.
+If this is possible in theory is discussed in the next paragraph.
+If it is exploitable in practice depends on the prices quoted on Deribit, liquidity and financing constraints.
+Since this question is of a more applied nature [^6] Complications can for example occur if as part of an initial issuance the company agrees to limits future issuance., it will be explored elsewhere.
+Can issuance of debt increase shareholder value in the Black-Scholes framework? This question is equivalent to finding a value $Q$ for which the expression
+$Y\,C_{BS}(B_{0},1\$\,Q/Y,\sigma,R,r,T)-B_{0}A$
+is positive. First replace, using the put-call parity, $C_{BS}(B_{0},1\$\,Q/Y,\sigma,R,r,T)$ to get
+$Y\,P_{BS}(B_{0},1\$\,Q/Y,\sigma,R,r,T)+YB_{0}e^{-rT}-1\$\,Qe^{-RT}-B_{0}A.$
+This is equivalent to
+$B_{0}Y(e^{-rT}-1)=(B_{0}A+XQ)(e^{-rT}-1),$
+which is equal to zero, for $r=0$, negative for $r>0$, and positive $r<0$.
+Therefore, zero-coupon bond issuance is for $r<0$ accretive for stockholders.
+At what stage the stock market recognises the gain, i.e.
+pre- or post-announcement, or maybe post-approval or -issuance, is debatable.
+Next, the analysis is extended to two issuances.
+We assume debt is issued in two tranches at $t=0$ with the amounts $Q_{1}$ and $Q_{2}$, and either priced sequentially with the existence of the second tranche not known when $X_{1}$, the fair value for the bond, is determined, or priced as one tranche of size $Q_{1}+Q_{2}$.
+Three cases are considered, for each there are three constraints and one inequality.
+The cases are named A,B & C, where A corresponds to one issuance of $Q_{1}$ units of debt, B corresponds to the issuance of $Q_{1}$ and $Q_{2}$ units of debt in sequence, but almost simultaneously, and C corresponds to one issuance of $Q_{1}+Q_{2}$ units of debt.
+I. The amount converted to Bitcoin is in the three scenarios
+$A:\,\,\,\,$ $AB_{0}+X_{1}Q_{1}=Y_{1}B_{0},$
+$B:\,\,\,\,$ $AB_{0}+X_{1}Q_{1}+X_{2}Q_{2}=Y_{2}B_{0},$
+$C:\,\,\,\,$ $AB_{0}+\overline{X}(Q_{1}+Q_{2})=\overline{Y}B_{0}.$
+II.The value of Put is in the three scenarios
+$A:\,\,\,\,$ $(e^{-RT}1\$\,-X_{1})Q_{1}=Y_{1}P\Big{(}B_{0},1\$\,\frac{Q_{1}}{Y_{1}},\sigma,R-r,T\Big{)},$
+$B:\,\,\,\,$ $(e^{-RT}1\$\,-X_{2})Q_{2}=\frac{Q_{2}}{Q_{1}+Q_{2}}Y_{2}P\Big{(}B_{0},1\$\,\frac{Q_{1}+Q_{2}}{Y_{2}},\sigma,R-r,T\Big{)},$
+$C:\,\,\,\,$ $(e^{-RT}1\$\,-\overline{X})(Q_{1}+Q_{2})=\overline{Y}P\Big{(}B_{0},1\$\,\frac{Q_{1}+Q_{2}}{\overline{Y}},\sigma,R-r,T\Big{)}.$
+III. Put-Call-Parity is in the three scenarios
+$A:\,\,\,\,$ $Y_{1}C\Big{(}B_{0},\frac{1\$\,Q_{1}}{Y_{1}},\sigma,r,T\Big{)}-Y_{1}P\Big{(}B_{0},\frac{1\$\,Q_{1}}{Y_{1}},\sigma,R,r,T\Big{)}=Y_{1}B_{0}e^{-rT}-e^{-RT}1\$\,Q_{1},$
+$B:\,\,\,\,$ $Y_{2}\Big{(}C\Big{(}B_{0},1\$\,\frac{Q_{1}+Q_{2}}{Y_{1}},\sigma,R,r,T\Big{)}-P\Big{(}B_{0},1\$\,\frac{Q_{1}+Q_{2}}{B_{1}},\sigma,R,r,T\Big{)}\Big{)}$
+$=Y_{2}B_{0}e^{-rT}-e^{-RT}1\$\,(Q_{1}+Q_{2}),$
+$C:\,\,\,\,$ $\overline{Y}\Big{(}C\Big{(}B_{0},1\$\,\frac{Q_{1}+Q_{2}}{\overline{Y}},\sigma,R,r,T\Big{)}-P\Big{(}B_{0},1\$\,\frac{Q_{1}+Q_{2}}{\overline{Y}},\sigma,R,r,T\Big{)}\Big{)}$
+$=\overline{Y}B_{0}e^{-rT}-e^{-RT}1\$\,(Q_{1}+Q_{2}).$
+IV. The Stock Value adjustment is therefore in the three scenarios
+$A:\,\,\,\,$ $Y_{1}C\Big{(}B_{0},1\$\,\frac{Q_{1}}{Y_{1}},\sigma,R,r,T\Big{)}-AB_{0},$
+$B:\,\,\,\,$ $Y_{2}C\Big{(}B_{0},1\$\,\frac{\,Q_{1}+\,Q_{2}}{Y_{2}},\sigma,R,r,T\Big{)}-AB_{0},$
+$C:\,\,\,\,$ $\overline{Y}C\Big{(}B_{0},1\$\,\frac{Q_{1}+Q_{2}}{\overline{Y}},\sigma,R,r,T\Big{)}-AB_{0}.$
+The first and third case are evaluated identically, and have been dealt with above, whilst for the second case the stock value change is
+$Q_{1}(X_{1}-X_{2}),$
+for $r=0$, and issuance of a zero coupon can at best be value neutral and otherwise destroys value. Next, $r\leq 0$ is considered. The stock value adjustment has now the form
+$(AB_{0}+Q_{2}X_{2})(e^{-rT}-1)+Q_{1}(e^{-rT}X_{1}-X_{2}),$
+which is positive for
+$-\frac{1}{T}\log\Big{(}\frac{(Q_{1}+Q_{2})X_{2}+AB_{0}}{AB_{0}+Q_{2}X_{2}+Q_{1}X_{1}}\Big{)}>r.$
+For small $r$ one has the expansion
+$Q_{1}(X_{1}-X_{2})-rT(AB_{0}+Q_{2}X_{2}+Q_{1}X_{1})+O(r^{2}).$
+Next a section on the asymmetric influence of uncertainty on optimal investment fractions.
+V The Kelly criterion in uncertain markets
 The one-dimensional Kelly criterion is in this section extended to uncertain winning probabilities.
@@
 $D_{\mathrm{KL}}\left(q\parallel p\right)=p\log\Big{(}\frac{p}{q}\Big{)}+(1-p)\log\Big{(}\frac{1-p}{1-q}\Big{)}$
-$=$ $\sum_{n=1}^{\infty}\Big{(}q+\epsilon\Big{)}(-1)^{n}\frac{\epsilon^{n}}{n\,\,q^{n}}-\Big{(}1-q-\epsilon\Big{)}\frac{\epsilon^{n}}{n\,\,(1-q)^{n}}$
-$=$ $\sum_{n=2}^{\infty}\frac{\epsilon^{n}}{n^{2}-n}\Big{(}\frac{(-1)^{n}}{q^{n-1}}-\frac{1}{(1-q)^{n-1}}\Big{)}=\sum_{n=2}^{\infty}\frac{\epsilon^{n}}{n^{2}-n}\Big{(}\frac{(-1)^{n}(1-q)^{n-1}-q^{n-1}}{q^{n-1}(1-q)^{n-1}}\Big{)}$
-$=$ $\frac{1}{2(1-q)q}\epsilon^{2}-\frac{1}{6}\frac{(1-q)^{2}+q^{2}}{(1-q)^{2}q^{2}}\epsilon^{3}+\frac{1}{12}\frac{(1-q)^{3}-q^{3}}{(1-q)^{3}q^{3}}\epsilon^{4}-\frac{1}{20}\frac{(1-q)^{4}+q^{4}}{(1-q)^{4}q^{4}}\epsilon^{5}+O(\epsilon^{6}).$
+$=$ $\sum_{n=1}^{\infty}\Big{(}q+\epsilon\Big{)}(-1)^{n+1}\frac{\epsilon^{n}}{n\,\,q^{n}}-\Big{(}1-q-\epsilon\Big{)}\frac{\epsilon^{n}}{n\,\,(1-q)^{n}}$
+$=$ $\sum_{n=2}^{\infty}\frac{\epsilon^{n}}{n^{2}-n}\Big{(}\frac{(-1)^{n+1}}{q^{n-1}}-\frac{1}{(1-q)^{n-1}}\Big{)}=\sum_{n=2}^{\infty}\frac{\epsilon^{n}}{n^{2}-n}\Big{(}\frac{(-1)^{n+1}(1-q)^{n-1}-q^{n-1}}{q^{n-1}(1-q)^{n-1}}\Big{)}$
+$=$ $\frac{1}{2(1-q)q}\epsilon^{2}-\frac{1}{6}\frac{(1-q)^{2}-q^{2}}{(1-q)^{2}q^{2}}\epsilon^{3}+\frac{1}{12}\frac{(1-q)^{3}+q^{3}}{(1-q)^{3}q^{3}}\epsilon^{4}-\frac{1}{20}\frac{(1-q)^{4}-q^{4}}{(1-q)^{4}q^{4}}\epsilon^{5}+O(\epsilon^{6}),$
+$=$ $\frac{1}{2(1-q)q}\epsilon^{2}+\frac{1}{6}\frac{2q-1}{(1-q)^{2}q^{2}}\epsilon^{3}+O(\epsilon^{4}).$
 If, as stated above, both $D_{\mathrm{KL}}\left(q\parallel p\right)$, equal to $\alpha$, and $q$ are fixed, then in lowest approximation, as one solves the equation up to a quadratic terms, one gets two solutions for $\epsilon$.
+One can extend the analysis to the cubic term without altering the result [^7] If one includes the cubic term and assumes reasonable values for $\alpha$ and $q$, one gets also in the relevant range two solutions $p_{1}\,\&\,p_{2}$ with $p_{1}\leq q\leq p_{2}$, then $|q-p_{1}|\geq|p_{2}-q|$ such that $p_{1}+p_{2}\leq 2q$.
+This would mean that up to cubic terms the average of the two individually ‘optimal’ fractions would be slightly lower than $q$.
+As a consequence, the optimal fraction is lower, i.e.
+$f=(2q-1)(1-\frac{2}{3}\alpha)+O(\alpha^{2})$, and also separately one can show that the growth rate is reduced by series in $\alpha$, i.e.
+$\alpha-\alpha^{2}\frac{(2q-1)^{2}}{18q(1-q)}+O(\alpha^{3})$.
+.
 Various cases can be considered.
@@
 with equal weights for both possible solutions of $p$.
-In the equal weighted case, due to the asymmetry associated with over- and under-investment, as described above, the optimal investment fraction is lower than $12q-$ 1.
+In the equal weighted case, due to the asymmetry associated with over- and under-investment, as described above, the optimal investment fraction is lower than 2q12q-1.
 To summarise, uncertainty about the right value for the winning probability leads to a lowering of the growth rate and the optimal fraction.
-An attractive formula for the adjusted investment fraction that captures this heuristically is $f^{*}\exp(-\lambda\alpha)$, where $f^{*}$ is the optimal uncertainty-free fraction $12q-$ 1, and $\lambda$ is a risk-aversion parameter.
+An attractive formula for the adjusted investment fraction that captures this heuristically is $f^{*}\exp(-\lambda\alpha)$, where $f^{*}$ is the optimal uncertainty-free fraction 2q12q-1, and $\lambda$ is a risk-aversion parameter.
 This result can be extended from KL divergence to the more general Bregman divergence by doing the relevant Taylor expansions for cases of interest like the Itakura–Saito distance.
 A similar ‘conservatism’ can be observed.
+The rationale for the fractional Kelly criterion, where one invests a fraction of the ‘optimal’ investment amount, can also be reevaluated in the light of the above result.
+A more conservative invest fractional investment is not due to excessive risk aversion, but instead due to uncertainty about the state of the world, i.e.
+in the simplest case the winning probability of a coin.
 The conclusion attempts again to shed light on treasury companies.
-V Conclusion
+VI Conclusion
 Crypto treasury companies with a rising ‘Bitcoin per share’ and mNAV above one seem to have discovered a financial perpetuum mobile and are eager to recruit followers, since copy-cats compete for the restricted supply of Bitcoins and in the process push up prices.
@@
 This could break the cycle, but, as the saying goes, ‘the market can stay irrational longer than investors can stay solvent’.
+This weakness is less evident for the strategy described in section IV based on debt issuance, which aims to generate instantaneous value for shareholders.
+Let’s recap: The idealised Bitcoin treasury company, which at the start solely holds crypto currency, issues a zero-coupon bond.
+The money raised is used to buy further Bitcoins.
+The bond is not risk-free, since Bitcoin falling below a threshold vis-a-vis dollar at bond maturity leads to bankruptcy.
+A riskless zero-coupon investment is the combination of the bond issued by the Bitcoin treasury company and a put, which indemnifies bond investors against losses in case of default.
+It was shown that the company value remains unaffected by an individual zero-coupon debt issuance of any size or maturity, independent of Bitcoin volatility, USD interest rate, and initial Bitcoins held by the company, if the Bitcoin interest rate is set to zero.
+Whilst issuance is accretive, if $r$ is negative.
+A more thorough investigation requires the study of more elaborate hybrid, i.e.
+equity plus debt, funding mechanisms which treasury companies are now employing.
+The analysis also ignored that one-sided activity leads to buying pressure on Bitcoin101010 Unlike the temporary price pressure from an index inclusion of an equity - which doesn’t normally affect fundamental value - this effect should be more pronounced for harder to evaluate cryptocurrencies, since the line between transitory and permanent market impact is more easily blurred.
+, and can, through ‘price-impact’, create market momentum.
+This effect and the resulting dynamic will be studied elsewhere together with general constructions involving leveraged positions in liquid assets.
 The history of finance is full of wondrous schemes.
@@
 In the second part of the paper, simple information theoretic ideas are employed to understand the influence of uncertainty on decision making, and an asymmetric effect on the optimal fraction for the binary-choice Kelly criterion was observed as the winning probability becomes uncertain.
-The paradox of high mNAV crypto treasury companies is captured by the White Queen in Lewis Carroll’s ‘Through the Looking Glass’, when she says, ‘[t]he rule is, jam to-morrow and jam yesterday – but never jam to-day’.
-  1. (1) J. Kelly, A new interpretation of information rate. Bell System Technical Journal 35(4): 917–926 (1956).
-  2.
+The paradox of high mNAV crypto treasury companies feverishly issuing equity is captured by the White Queen in Lewis Carroll’s ‘Through the Looking Glass’, when she says, ‘[t]he rule is, jam to-morrow and jam yesterday – but never jam to-day’.
+  1. (1) P. Bergault, S. Bieber, O. Gueant, and W. Zhang. “Cryptocurrencies and Interest Rates: Inferring Yield Curves in a Bondless Market.” Available at SSRN 5321931, & arXiv:2509.03964 (2025).
+  2. (2) D.C. Brody, L.P. Hughston, & B.K. Meister. “Theory of cryptocurrency interest rates.” SIAM Journal on Financial Mathematics 11, no. 1 (2020): 148-168.
+  3. (3) J. Kelly, “A new interpretation of information rate.” Bell System Technical Journal 35(4): 917–926 (1956).
+  4.
 (4) Y.
@@
   8. (8) B. K. Meister, “Application of the Kelly Criterion to Prediction Markets ”, arXiv:2412.14144. (2024).
-  7. (7) S. Nakamoto, ‘Bitcoin: A Peer-to-Peer Electronic Cash System’ (2008).
+  9. (9) S. Nakamoto, “Bitcoin: A Peer-to-Peer Electronic Cash System” (2008).